### v1

Abstract
We initiate the study of multipacking problems for geometric point sets with respect to their Euclidean distances. We consider a set of $n$ points $P$ and define $N_{s}[v]$ as the subset of $P$ that includes the $s$ nearest points of $v\in P$ and the point $v$ itself. We assume that the $s$-th neighbor of each point is unique, for every $s\in\{0,1,2,\dots,n-1\}$. For a natural number $r\leq n$, an $r$-multipacking is a set $M\subseteq P$ such that for each point $v\in P$ and for every integer $1\leq s\leq r$, $|N_{s}[v]\cap M|\leq(s+1)/2$. The $r$-multipacking number of $P$ is the maximum cardinality of an $r$-multipacking of $P$ and is denoted by $\operatorname{mp}_{r}(P)$. For $r=n-1$, an $r$-multipacking is called a multipacking and $r$-multipacking number is called as multipacking number. We study the problem of computing a maximum $r$-multipacking for point sets in $\mathbb{R}^{2}$. We show that a maximum $1$-multipacking can be computed in polynomial time but computing a maximum $2$-multipacking is NP complete. Further, we provide approximation and parameterized solutions to the $2$-multipacking problem.
Keywords:
Geometric packing Nearest neighbor graph Multipacking.
1 Introduction
The problem of locating facilities, both desirable and undesirable, is of utmost importance due to its huge number of applications in the real world [22]. Allocation of restricted resources within an arena has been extensively studied across various models and methodologies since the 1980s. Drezner and Wesolowsky did a "maximin" problem where a certain facility must serve a given set of users but must be placed in such a manner that its minimum distance from the users is maximized (see [11, 12]). Rodriguez et al. [20] gave a generic model for the obnoxious facility location problem inside a given polygonal region. Rakas et al. [19] devised a multi-objective model to determine the optimal locations for undesirable facilities. The book by Daskin [9] contains a nice collection of problems, models, and algorithms regarding obnoxious facility location problems in the domain of supply chain networks. In this paper, we initiate a novel study of the Multipacking [21] problem in Euclidean space, which resembles the obnoxious facility location problem in computational geometry.
Given a planar point set $P\subseteq\mathbb{R}^{2}$ of size $n$, the neighborhood of size $r$ of a point $v\in P$ is the subset of $P$ that consists of the $r$ nearest points of $v$ and $v$ itself. This subset is denoted by $N_{r}[v]$. A natural and realistic assumption here is the distances between all pairs of points are distinct. Thus the $r$-th neighbor of every point is unique. This assumption is easy to achieve through some form of controlled perturbation (see [16]). For a natural number $r\leq n$, an $r$-multipacking is a subset $M$ of $P$, such that for each point $v\in P$ and for every integer $1\leq s\leq r$, $|N_{s}[v]\cap M|\leq(s+1)/2$, i.e. $M$ contains at most half of the number of elements of $N_{s}[v]$. The $r$-multipacking number of $P$ is the maximum cardinality of an $r$-multipacking of $P$ and is denoted by $\operatorname{mp}_{r}(P)$.
An $(n-1)$-multipacking $M$ of $P$ is referred to as multipacking of $P$. The multipacking number of $P$, denoted by $\operatorname{mp}(P)$, refers to the cardinality of a maximum multipacking of $P$. Computing maximum multipacking involves selecting the maximum number of points from the given point set satisfying the constraints of multipacking. This variant of facility location problem is interesting due to their wide application in operation research [20], resource allocation [18], VLSI design [1], network planning [6, 19], and clustering [15].
To the best of our knowledge, no significant work on maximum multipacking has been done on planar point sets concerning their Euclidean distances. Whereas the literature contains numerous intriguing findings for this problem in graph theory. Multipacking was first discussed by Teshima [21]. Other related works can be found in [2, 7, 8, 17]. However, as of now, no polynomial-time algorithm is known for finding a maximum multipacking of general graphs, the problem is also not known to be NP-hard. Nonetheless, polynomial-time algorithms exist for trees and strongly chordal graphs [5]. For further references on algorithmic results regarding multipacking problem on graphs, we refer to the paper [14].
Further, we define the function $\operatorname{MP}_{r}(t)$ to be the smallest number such that any point set $P\subset\mathbb{R}^{2}$ of size $\operatorname{MP}_{r}(t)$ admits a maximum $r$-multipacking of size exactly $t$. We denote $\operatorname{MP}_{n-1}(t)$ as $\operatorname{MP}(t)$, i.e. $\operatorname{MP}(t)$ is the smallest number such that any point set $P\subset\mathbb{R}^{2}$ of size $\operatorname{MP}(t)$ admits a maximum multipacking of size exactly $t$. It can be followed that $\operatorname{MP}(1)=1$ from the definition of multipacking. No general solution has been found for this problem yet.
Our Contribution: In Section 2, we provide an algorithm to find a maximum multipacking of a point set on a line and we prove some bounds on the multipacking number on a line. In Section 3, we prove that $\operatorname{MP}(2)=6$ using some geometric observations. Until now, there is no known polynomial-time algorithm to find a maximum multipacking for a given point set $P\subset\mathbb{R}^{2}$, and the problem is also not known to be NP-hard. In this paper we make an important step towards solving the problem by studying the $1$-multipacking and $2$-multipacking in $\mathbb{R}^{2}$. In Section 4, we considered the NNG (Nearest-Neighbor-Graph) $G$ of $P\subset\mathbb{R}^{2}$. Using the observation that the maximum independent sets of $G$ are the maximum $1$-multipackings of $P$, we conclude that a maximum $1$-multipacking can be computed in polynomial time. However the $2$-multipacking problem in $\mathbb{R}^{2}$ is much more complicated. In Section 5, we study the hardness of the $2$-multipacking problem in $\mathbb{R}^{2}$. We prove that the problem is NP-complete by reducing the well-known NP-complete problem Planar Rectilinear Monotone 3-SAT [10] to the $2$-multipacking problem. Moreover, we provide approximation and parameterized solutions to this problem.
2 Maximum multipacking in $\mathbb{R}^{1}$
2.1 Bound of maximum multipacking in $\mathbb{R}^{1}$
We start by bounding the multipacking number of a point set in the following lemma.
Lemma 1
Let $P$ be a point set on $\mathbb{R}^{1}$. Then $\lfloor\frac{n}{3}\rfloor\leq\operatorname{mp}(P)\leq\lfloor\frac{n}{2}\rfloor$.
Proof
In order to show the lower bound, we show that for any point set on $\mathbb{R}^{1}$ of size $n$ we can always find a multipacking of size $\lfloor\frac{n}{3}\rfloor$. Let $P=\{p_{1},p_{2},\dots,p_{n}\}$ be $n$ points on $\mathbb{R}^{1}$ sorted with respect to the ascending order of their $x$-coordinates. Construct a set $M$ as such, $M=\{p_{1},p_{4},p_{7},\dots,p_{3k+1}\}$ (where $k\in\mathbb{N}$ and $13k+1\leq n<3(k+1)+$ 1) is a multipacking of $P$. Note that $M$ is a multipacking since for any $11\leq s\leq n-$ 1 and $p_{i}\in P$, $N_{s}[p_{i}]$ contains at most $\lfloor\frac{s+1}{2}\rfloor$ points from $M$. This shows the lower bound of the lemma. The upper bound follows from the definition of multipacking. $\square$
Tight bound example for the upper bound: For the upper bound of Lemma 1, we consider the point set $P=\{p_{1},\ldots,p_{n}\}$ where,
$p_{i}=\begin{cases}0&\text{if $i=0$ }\\ \frac{4}{3}(2^{i-1}-1)-\frac{i-1}{2}&\text{if $i$ is odd}\\ \frac{4}{3}(2^{i}-1)-\frac{i}{2}-2^{i-1}+1&\text{if $i$ is even, $i>0$}\end{cases}$
Note that $p_{i}$’s in $P$ are in ascending order. From definition of this example, $p_{2k+2}-p_{2k+1}>p_{2k+1}-p_{1}$ and $p_{2k+2}-p_{2k+1}>p_{2k+3}-p_{2k+2}$ for each $k$. From this, we conclude that the set $M=\{p_{1},p_{4},p_{6},p_{8},\ldots\}$ is a maximum multipacking and $|M|=\lfloor\frac{n}{2}\rfloor$.
Tight bound example for the lower bound: For the lower bound of Lemma 1, we consider the point set $P=\{p_{1},\ldots,p_{n}\}$ where $p_{i}=2^{i}$.
Note that $p_{i-1}-p_{1}<p_{i}-p_{i-1}$ holds for each $p_{i}$ in $P$. From this, we conclude that the set $M=\{p_{1},p_{4},p_{7},\ldots\}$ is a maximum multipacking with $|M|=\lfloor\frac{n}{3}\rfloor$.
Combining these arguments together with Lemma 1 we have the following theorem.
Theorem 2.1 ()
Let $P$ be a point set on $\mathbb{R}^{1}$. Then $\lfloor\frac{n}{3}\rfloor\leq\operatorname{mp}(P)\leq\lfloor\frac{n}{2}\rfloor$. Both the bounds are tight.
2.2 Algorithm to find a maximum multipacking in $\mathbb{R}^{1}$
We devise algorithm 2 that checks possible violations for every member of $P$ and selects the eligible candidates as multipacking accordingly. It takes a point set $P\subset\mathbb{R}^{1}$ sorted in ascending order with respect to their $x$-coordinates as input and outputs a maximum multipacking $M$ where $M\subseteq P$.
Algorithm 1 ($CMP_{r}(P,M)$) is designed to check the eligibility of a member of $P$ as a point in multipacking and executed as a subroutine in each iteration of algorithm 2. $CMP_{r}(P,M)$(Algorithm 1) takes a point set $M\subseteq P$ as input and outputs $1$ if $M$ is a multipacking and 00 if not. Given a point set $P=\{p_{1},p_{2},\ldots,p_{n}\}$, at the $i^{\text{th}}$ iteration, $p_{i}$ is included in $M$ and $CMP_{r}(P,M)$ is called as a subroutine. Depending on the eligibility check done by this subroutine $M$ is formed. The algorithm terminates when $P$ is exhausted and returns $M$ as output. The correctness of the algorithm is presented in the appendix [proof of Theorem 2.2].
Input: A point set $P$ in $\mathbb{R}$ and $M\subseteq P$ and the value of $r$.
Output: if $M$ is a valid multipacking on $P$ then return $1$else return 00
for $p_{i}$ do
       for $s=1\rightarrow r$ do
             if $|N_{s}[p_{i}]\cap M|\leq(s+1)/2$ then
                  continue;
            else
                  return 00;
             end if
       end for
end for
return $1$;
Algorithm 1 Check r-multipacking - $CMP_{r}(P,M)$
Input: A point set $P$ in $\mathbb{R}$ and the value of $r$.
Output: A maximum multipacking $M$ of $P$.
$M=\phi$;
$i=1$;
for $p_{i}$ do
       $M=M\cup\{p_{i}\}$;
       if $CMP_{r}(P,M)==0$ then
            $M=M\setminus\{p_{i}\}$;
       end if
      $i++$;
end for
return $M$;
Algorithm 2 Maximum $r$-multipacking
Theorem 2.2 ()
For any $11\leq r\leq n-$ 1, a maximum $r$-multipacking for a given point set $P\subset\mathbb{R}^{1}$ with $|P|=n$ can be computed in $O(n^{2}r)$ time.
It is worth mentioning that, computing the maximum multipacking set of a planar point set relates with the notion of independent set (which we show later in the paper) which becomes NP-hard in general. Thus it is important to study a bound for the multipacking number in $\mathbb{R}^{2}$. Next, we address the bound on the size of $P$ to achieve a multipacking of a desired size irrespective of the distribution of a set of points in $\mathbb{R}^{2}$.
3 Minimum number of points to achieve a desired multipacking size in $\mathbb{R}^{2}$
In this section, we study the $\operatorname{MP}$ function. Recall the definition of the same. $\operatorname{MP}_{r}(t)$ is the smallest number such that any point set $P\subset\mathbb{R}^{2}$ with $|P|=\operatorname{MP}_{r}(t)$ admits an $r$-multipacking of size exactly $t$. As stated earlier $\operatorname{MP}_{n-1}(t)$ is denoted by $\operatorname{MP}(t)$ and $\operatorname{MP}(1)=1$. Now our goal is to find the value of $\operatorname{MP}(2)$, which we state in the following theorem.
Theorem 3.1
$\operatorname{MP}(2)=6$.
Figure 1: $6$ points with $\operatorname{MP}(2)\leq 6$.
Figure 2: $5$ points with $\operatorname{MP}(2)>5$.
Proof
We prove this using contradiction. Suppose $\operatorname{MP}(2)>6$. That means there is a point set $P=\{a_{1},a_{2},\dots,a_{6}\}\subset\mathbb{R}^{2}$ such that $\operatorname{mp}(P)=1$. This implies, for each $i,j(i\neq j)$ there exists $k$ such that $\{a_{i},a_{j}\}\subset N_{2}[a_{k}]$, otherwise $\{a_{i},a_{j}\}$ can form a multipacking of size $2$ which is not allowed. If $a_{i}$ and $a_{j}$ are the $1^{\text{st}}$ and $2^{\text{nd}}$ neighbor of $a_{k}$ respectively, we denote $C_{2}[a_{k}]$ to be the circle with center at $a_{k}$ passing through $a_{j}$. Without loss of generality we can assume that $C_{2}[a_{1}]$ is the circle with the largest radius among all the circles $C_{2}[a_{k}]$, $k\in\{1,2,\dots,6\}$ and moreover $a_{2}$ and $a_{3}$ are the $1^{\text{st}}$ and $2^{\text{nd}}$ neighbor of $a_{1}$ (See Fig. 2). Observe that, $a_{1}\notin N_{2}[a_{k}]$ for each $k\in\{4,5,6\}$, since $C_{2}[a_{1}]$ is the circle with largest radius. Therefore, each pair among the three pairs $\{a_{1},a_{4}\}$, $\{a_{1},a_{5}\}$ and $\{a_{1},a_{6}\}$ is a subset of either $N_{2}[a_{2}]$ or $N_{2}[a_{3}]$. By Pigeonhole Principle, there are at least two pairs among the three pairs $\{a_{1},a_{4}\}$, $\{a_{1},a_{5}\}$ and $\{a_{1},a_{6}\}$ which are subsets of either $N_{2}[a_{2}]$ or $N_{2}[a_{3}]$, but both $N_{2}[a_{2}]$ and $N_{2}[a_{3}]$ have size $3$ which is a contradiction of $\operatorname{mp}(P)=1$. Therefore, $\operatorname{MP}(2)\leq 6$.
Furthermore, we have point sets of sizes $3,4$ and $5$ (See Fig. 2) that have multipacking number $1$. We depict the point set $P$ in Fig. 2 forming a convex pentagon. Every point $p_{i}\in P$ has $p_{i-1}$ and $p_{i+1(mod~{}5)}$ as its two nearest neighbors. As a result no two points can be present in the multipacking set of $P$. The correctness of the figure can be followed from the fact that each circle (marked with different colours) containing a point $p_{i}\in P$ as center and the second nearest neighbor of $p_{i}$ on its boundary, has only one other point (namely the first nearest neighbor of $p_{i}$) in the interior. This example shows $\operatorname{MP}(2)>5$ implying, $\operatorname{MP}(2)=6$. $\square$
4 $1$-multipacking in $\mathbb{R}^{2}$
We note that algorithm 2 in subsection 2.2 can not be adapted for a point set $P\subseteq\mathbb{R}^{2}$. Since we can not ensure the sorted order of the points with respect to their $x$ or $y$ coordinate, the maximum cardinality of the output can not be guaranteed. Thus it is interesting to address $r$-multipacking while $r<n-1$.
The definition of $r$-multipacking for $r=1$ coincides with the definition of independent set in case of graphs and it remains the same when we consider the maximum independent set of the NNG (Nearest-Neighbor-Graph) of $P$. The NNG of $P$ is a forest when considered as an undirected graph with $2$-cycles at the leaves. Then, a maximum $1$-multipacking for $P$ can be computed in polynomial time by computing the maximum independent set of NNG of $P$ [13].
We write this observation as the following lemma.
Lemma 2
Let $P$ be a point set on the $2$D plane and the $NNG$ of $P$ be $G$. The maximum independent sets of $G$ are the maximum $1$-multipackings of $P$.
5 $2$-multipacking in $\mathbb{R}^{2}$
Next we study the $2$-multipacking problem. We are interested in showing the hardness of finding a $2$-multipacking of a given set of points $P$ in the plane.
We construct a graph $G_{P}=(V,E)$ from the given point set $P$. Here the set of vertices $V=P$. For every vertex $v\in V$, we introduce three edges $vu_{1},vu_{2},$ and $u_{1}u_{2}$ where $u_{1},u_{2}\in V$ are the first and second neighbors of $v$ respectively. The following lemma shows the relation between the independent set of $G_{P}$ and a $2$-multipacking of $P$.
Lemma 3
A set $M$ is a $2$-multipacking of $P$ if and only if the vertices in the graph $G_{P}$ corresponding to the members of $M$ form an independent set in $G_{P}$.
Proof
It follows from the definition of $2$-multipacking that for every tuple $(v,u_{1},$ $u_{2})$, as described above, at most one can be chosen as a member of $M$. Since each edge of $G_{P}$ is incident on a pair of vertices appearing in a tuple, there is no edge between two members of $M$ appearing in $G_{P}$. Thus the vertices corresponding to the members of $M$ form an independent set of $G_{P}$. The converse follows from a similar argument. $\square$
Thus, finding a maximum independent set of $G_{P}$ is equivalent to finding a maximum $2$-multipacking of $P$. In the rest of this section, we show that finding a maximum independent set of $G_{P}$ is NP-complete, and we devise approximation and parameterized solutions to this problem.
5.1 NP-completeness
We show that finding the maximum independent set in the graph $G_{P}$ is NP-complete by reducing the well-known NP-complete problem Planar Rectilinear Monotone (PRM) 3-SAT [10] to this problem. This variant of the boolean formula contains three literals in each clause and all of them are either positive or negative. Moreover, a planar graph can be constructed for the formula such that a set of axis parallel rectangles represent the variables and the clauses with the following properties. The rectangles representing the variables lie on the $x$-axis. Whereas rectangles representing the clauses that contain the positive (negative) literals lie above (below) the $x$-axis. Additionally, the rectangles for the clauses can be connected with vertical lines to the rectangles for the variables that appear in the clauses.
Given a PRM 3-SAT formula $\phi$ with $t(>1)$ variables and $m(>1)$ clauses, we construct a planar point set $P_{\phi}$ such that $\phi$ is satisfiable if and only if $G_{P_{\phi}}$ has a maximum independent set of a given size. First, we construct a point set for the variable gadget for each variable that has only two subsets as the maximum independent set in $G_{P_{\phi}}$. Then we construct another point set as clause gadgets for each clause in $\phi$ such that the clause gadget can incorporate only one vertex in the independent set only if at least one variable gadget has the desired set chosen as the maximum independent set. The clause and variable gadgets are connected via connection gadgets which are two point sets referred to as translation and rotation gadgets. We describe the construction of the point sets formally below.
Figure 3: A variable gadget in $G_{P_{\phi}}$.
Figure 4: A clause gadget and its connection with the variable gadgets.
Variable gadget. We construct a point set for each variable $x_{i}$ in $\phi$ consisting of $p=6(m-1)+2$ points. A variable gadget is depicted in Figure 4 with $p=8$. Each point is considered as the vertex in the graph $G_{P_{\phi}}$. The leftmost point is $v_{i}^{1}$ having $v_{i}^{2}$ and $v_{i}^{3}$ as its first and second neighbor respectively. Thus $v_{i}^{1}$ introduces three edges $v_{i}^{1}v_{i}^{2}$, $v_{i}^{1}v_{i}^{3}$, and $v_{i}^{2}v_{i}^{3}$ in $G_{P_{\phi}}$. The next point from the right is $v_{i}^{2}$ which has $v_{i}^{1}$ and $v_{i}^{3}$ has its first and second neighbor. Since the edges corresponding to the adjacency of $v_{i}^{2}$ are already drawn, it does not introduce any new edges. Then the third point $v_{i}^{3}$ introduces two new edges namely $v_{i}^{3}v_{i}^{4}$ and $v_{i}^{2}v_{i}^{4}$. Continuing in a similar fashion $v_{i}^{p-1}$ introduces two edges $v_{i}^{p-1}v_{i}^{p}$ and $v_{i}^{6}v_{i}^{p}$. Finally, $v_{i}^{p}$ does not introduce any new edges having $v_{i}^{p-1}$ and $v_{i}^{6}$ as its first and second neighbors respectively. The second neighborhoods of the points are indicated with colored (referring to the online version in case of gray-scale documents) circles in the figure. The leftmost red circle contains the first and second neighbor of $v_{i}^{1}$ in Figure 4. The following observation holds from the structure of the variable gadget in $G_{P_{\phi}}$.
Observation 1
The maximum independent set in a variable gadget with $26(m-1)+$ 2 vertices, as described above, is of size $12m-$ 1. There are exactly two options for selecting these two sets, either the sets consisting of the vertices $V_{i}^{T}=\{v_{i}^{1},v_{i}^{4},\dots,v_{i}^{p-1}\}$ or the sets consisting of the vertices $V_{i}^{F}=\{v_{i}^{2},v_{i}^{5},\dots,v_{i}^{p}\}$. Moreover, we can place the variable gadgets in such a way that the vertices $v_{i}^{2},v_{i}^{4},\dots,v_{i}^{p-1}$ lie above the $x$-axis and the other vertices lie below the $x$-axis.
We interpret the assignments of the variable $x_{i}$ to be true (false resp.) if $V_{i}^{T}$ ($V_{i}^{F}$) is chosen as the maximum independent set in the variable gadget corresponding to $x_{i}$. Now we describe the structure of a clause gadget.
Clause gadget. A clause gadget consists of a point set such that the corresponding graph contains $7$ triangles. We describe the construction for a positive clause $C=(x_{i}\lor x_{j}\lor x_{k})$ and the construction is analogous for the other clauses. The central triangle of the gadget in $G_{P_{\phi}}$ consists of three vertices $c_{x_{i}},c_{x_{j}}$, and $c_{x_{k}}$ as shown (with pink shade) in Figure 4. The triangle is introduced by placing $c_{x_{j}}$ and $c_{x_{k}}$ as the first and second neighbor of $c_{x_{i}}$ respectively. Then we draw three other vertices for each variable appearing in the clause. We describe the position of them for $x_{i}$ and the others are analogous. $p_{i},q_{i}$ and $r_{i}$ are the three vertices placed near $c_{x_{i}}$ such that $q_{i}$ has $c_{x_{i}}$ as its first neighbor and $p_{i}$ as its second neighbor. This introduces $\triangle q_{i}p_{i}c_{x_{i}}$ in $G_{P_{\phi}}$. Then $p_{i}$ has $q_{i}$ and $r_{i}$ as its first and second neighbor respectively introducing $\triangle p_{i}q_{i}r_{i}$ in $G_{P_{\phi}}$. We adjust the neighborhood of the other vertices are adjusted in such a way that no other triangle is introduced in $G_{P_{\phi}}$. For example, the first and second neighbors of $r_{i}$ are $q_{i}$ and $p_{i}$ respectively and the first and second neighbors of $c_{x_{j}}$ are $c_{x_{i}}$ and $c_{x_{k}}$ respectively. Considering the position of the points in the clause gadget following observation holds for the maximum independent set in the induced graph with the points in the clause gadget as the vertices.
Observation 2
The maximum independent set in a clause gadget as described above has size $4$ if and only if at least one of the three vertices $r_{i},r_{j}$, and $r_{k}$ is chosen in the independent set.
We consider an ordering of the clauses from left to right in the planar embedding of $\phi$. Let $C$ be the $\xi$-th clause that is connected with the variable $x_{i}$ from left to right. We translate each of the triangles $\triangle p_{i}q_{i}r_{i}$, $\triangle p_{j}q_{j}r_{j}$ and $\triangle p_{k}q_{k}r_{k}$ near to the corresponding variable gadgets to connect them. We place three points $p^{\prime}_{i},q^{\prime}_{i}$,and $r^{\prime}_{i}$ near the point $v_{i}^{2+6(\xi-1)}$ such that $r^{\prime}_{i}$ has $v_{i}^{2+6(\xi-1)}$ and $p^{\prime}_{i}$ as its first and second neighbour. We call the vertex $v_{i}^{2+6(\xi-1)}$ as the connecting vertex of $C$ and $x_{j}$. Then we construct a point set such that the maximum independent set induced by these points as vertices in $G_{P_{\phi}}$ has the following property. $p_{i}$($q_{i}$ or $r_{i}$ resp.) can be chosen in the maximum independent set if and only if $p^{\prime}_{i}$ ($q^{\prime}_{i}$ or $r^{\prime}_{i}$) is also chosen in the maximum independent set. We call this property as the connection of the clause to the variable gadgets. Once the connections to every clause and its variables are complete we have the following observation.
Observation 3
The vertex $c_{x_{i}}$ can be chosen in an independent set in $G_{P_{\phi}}$ only if the maximum independent set of the variable gadget corresponding to $x_{i}$ is chosen to be $V_{i}^{T}$.
The remaining work is to achieve the connection from $C$ to $x_{i},x_{j}$, and $x_{k}$. We use two point sets (gadgets) to perform this move. The first one is a translating gadget and the second one is a rotating gadget depicted in Figure 9 in appendix. The following observation can be verified from the structure of the gadgets.
Observation 4
Both the gadgets contain $6$ points inside them and the size of the maximum independent set is $2$.
The translating gadget starts with a copy of a triangle and propagates the choice of the vertex chosen in an independent set in the underlying graph. The rotating gadget does the same thing but the final triangle in this gadget is a copy of the initial triangle in the gadget but rotated at a right angle. Thus we can use mirror-reflections of the rotating gadget depicted in the figure to achieve the rotating angle to be $180^{\circ}$ and $270^{\circ}$. Together we call them the connecting gadgets.
Now the following observation is evident from the structure of the clause gadget and its connection with the variables.
Observation 5
Let $C=(x_{i}\lor x_{j}\lor x_{k})$ one clause that is connected with its variables in $G_{P_{\phi}}$ with $\nu$ connecting gadgets then the maximum independent set for the point set has size $3(2m+1)+(\nu+4)$ if and only if at least one of the three variables is assigned to be true.
Now we adjust our drawing of the whole point set for $G_{P_{\phi}}$ such that the size of the independent set is a fixed number depending on $\phi$. We place each clause gadget $C=(x_{i}\lor x_{j}\lor x_{k})$ in such a way that the connecting gadgets for $x_{j}$ do not contain a rotating gadget for the connection of $c_{x_{j}}$ to its connecting vertex in $x_{j}$. We consider the fact that the nearest clause gadget from the $x$-axis can be connected with its variables by two rotating gadgets for the first and third variables and one translating gadget for the middle variable appearing in the clause. Then we compute the maximum number of connecting gadgets required to connect a clause with the variables. Let $l_{v}$ and $w_{v}$ denote the length and width of the rectangles representing the variables in the planar drawing of $\phi$. We place our variable gadgets inside the rectangles. So $l_{v}=O(m)$ as we have placed $p=6(m-1)+2$ points in each variable gadget. Since there are $t$ variables and every clause rectangle can be connected with the variable rectangles using vertical lines, then the maximum length of a rectangle representing a clause could be $tl_{v}+\varepsilon$ for some small constant $\varepsilon$. On the other hand, if $w_{c}$ and $h_{c}^{0}$ denote the width of the clause-rectangle and the height (distance from $x$-axis) of the nearest clause-rectangle from the $x$-axis, then the distance of the farthest clause rectangle from the $x$-axis is $mw_{c}+h_{c}^{0}$. Then the maximum number of connecting gadgets required for connecting a clause gadget to its variables is at most $3mw_{c}+2tl_{v}$. Thus the clause gadget together with the translating and rotating gadget consists of $126(3mw_{c}+2tl_{v})+$ 12 points. And thus contributing a $42(3mw_{c}+2tl_{v})+$ 4 size maximum independent set. We draw our gadget in such a way that $w_{c}$ contains $6$ and $l_{v}$ contains $p$ points. Thus, every clause contains 6($126(18m+2tp)+$ 12 points together with its connecting gadgets.
To achieve this drawing, We expand the planar embedding of $\phi$ by a scalar $\alpha=(10tm)^{2}$. The size of the gadgets remains the same but the space between them increases such that there is a distance of at least $\alpha$ between any two connecting gadgets to accommodate the total number of points. The convolution to accommodate the extra points for the connection is depicted in Figure 7. A lower-level diagram for the embedding of such convolution with translating and rotating gadgets is depicted in Figure 9 in appendix. Then we claim the following lemma.
Lemma 4
Given a PRM-3SAT formula $\phi$ we can construct a planar point set $P_{\phi}$ such that $G_{P_{\phi}}$ has a maximum independent set of size $t(m+1)+m(2(18m+2tp)+4)$ if and only if $\phi$ is satisfiable.
Since it is possible to check the correctness of a solution for an independent set we can conclude the following theorem.
Theorem 5.1 ()
Computing a maximum $2$-multipacking of a given set of $n$ points $P$ in $\mathbb{R}^{2}$ is NP-complete.
5.2 An Approximation and a Parameterised Algorithm
In this subsection, we proved that the maximum degree of $G_{P}$ is bounded. The proof of the lemma stated below is given in the appendix.
Lemma 5
The maximum degree of the graph $G_{P}$ is at most $17$.
Using this we have shown that the $2$-multipacking of a given set of points $P$ in the plane can be approximated in polynomial time. We use existing techniques for computing the maximum independent set for bounded degree graphs.
The problem of finding the Maximum Independent Set in a bounded degree graph is denoted by MAX IS-$\Delta$ where the maximum node degree of the graph is bounded above by $\Delta$. The following result states the approximation bound of the MAX IS-$\Delta$ problem.
Theorem 5.2 (Berman and Fujito [3])
MAX IS-$\Delta$ can be approximated in polynomial time within a ratio arbitrarily close to $(\Delta+3)/5$ for all $\Delta\geq 2$.
By Theorem 5.2 and Lemma 5, we have the following result stated earlier.
Theorem 5.3 ()
A maximum $2$-multipacking of a given set of points $P$ in $\mathbb{R}^{2}$ can be approximated in polynomial time within a ratio arbitrarily close to $4$.
In 2019, Bonnet et al. have shown the following:
Theorem 5.4 (Bonnet et al.[4])
For graphs of $n$ vertices with degree bounded by $\Delta$, we can compute an independent set of size $k$ in time $(\Delta+1)^{k}n^{O(1)}$, if it exists.
Lemma 5 and Theorem 5.4 yield the following result stated earlier.
Theorem 5.5 ()
A $2$-multipacking of size $k$ of a given set of $n$ points in $\mathbb{R}^{2}$ can be computed in time $18^{k}n^{O(1)}$, if it exists.
6 Conclusion
The paper introduces a novel modeling of the maximum multipacking problem on the geometric domain concerning the Euclidean distance for a point set. An intriguing open topic in this area is addressing the difficulty of computing a maximum multipacking in general. We know neither the algorithmic difficulties nor the limitations as of yet for the $(n-1)$-multipacking problem.
The definition in this paper is motivated by problems in obnoxious facility location and multipacking in graph theory. In the definition given in this paper, one can introduce a fraction $k\in[0,1]$ and generalise the given condition with respect to this fraction. For $k\in[0,1]$, the condition on set $M$ can be given as such : $|N_{s}[v]\cap M|\leq k(s+1)$ for each point $v\in P$.
For different values of $k$, the bounds on the multipacking number as well as the algorithmic difficulties will differ significantly.